OpenStudy (anonymous):

What is the equation of the line that passes through the points (-2, 1) and (1, 10)?

OpenStudy (anonymous):

3x - y = -7 3x - y = -5 3x - y = 5 x + 3y = -5

OpenStudy (anonymous):

your choices are not in slope form, but here is the correct equivalent.